If I've been given the perimeter of a rectangle which is 360 and given the length as 8x-9 how can I find my width (x) so that I have the dimensions of the rectangle.
Answer by amarjeeth123(569)

Perimeter=2(length+width)
2(8x-9+x)=360
2(9x-9)=360
18(x-1)=360
x-1=20
x=21
The width of the rectangle is 21.
The length of the rectangle is 159.